Rachels family drove 984 miles in 2 days. If the ratio of the distance covered each day was 20 to 21, find the number of miles they drove each day.
Answer by ankor@dixie-net.com(22740)

Rachels family drove 984 miles in 2 days.
If the ratio of the distance covered each day was 20 to 21,
find the number of miles they drove each day.
:
let x = the multiplier
then
20x = no. of miles on day 1
and
21x = no. of miles on day 2
:
20x + 21x = 984
41x = 984
x = 984/41
x = 24 is the multiplier
:
20*24 = 480 on day 1
21*24 = 504 on day 2
----------------------
total: 984 mi for two days